### v1

# HiMat: DiT-based Ultra-High Resolution SVBRDF Generation

Abstract.
Creating highly detailed SVBRDFs is essential for 3D content creation. The rise of high-resolution text-to-image generative models, based on diffusion transformers (DiT), suggests an opportunity to finetune them for this task. However, retargeting the models to produce multiple aligned SVBRDF maps instead of just RGB images, while achieving high efficiency and ensuring consistency across different maps, remains a challenge. In this paper, we introduce HiMat: a memory- and computation-efficient diffusion-based framework capable of generating native 4K-resolution SVBRDFs. A key challenge we address is maintaining consistency across different maps in a lightweight manner, without relying on training new VAEs or significantly altering the DiT backbone (which would damage its prior capabilities). To tackle this, we introduce the CrossStitch module, a lightweight convolutional module that captures inter-map dependencies through localized operations. Its weights are initialized such that the DiT backbone operation is unchanged before finetuning starts. HiMat enables generation with strong structural coherence and high-frequency details. Results with a large set of text prompts demonstrate the effectiveness of our approach for 4K SVBRDF generation. Further experiments suggest generalization to tasks such as intrinsic decomposition.
material generation, SVBRDF, generative models
submissionid: 1275journal: TOGccs: Computing methodologies Rendering
Figure 1. We present HiMat, a diffusion-based model for ultra-high-resolution ($4096\times 4096$) SVBRDF material generation from text prompts. Our approach achieves native 4K-resolution with both memory and time efficiency, while preserving high-frequency details crucial for mesostructural components such as normal and height maps. See the materials generated by our method for each individual chess piece, with zoom-in views that highlight the preserved fine-scale details and texture fidelity.
1. Introduction
Generating detailed materials in the form of high-resolution (e.g., $4096\times 4096$) SVBRDFs would be highly desirable for enhancing realism in 3D content creation, to faithfully capture fine surface detail and support close-up rendering. While some text-to-image generative models (Xie et al., 2025a) have achieved this for 3-channel RGB image generation, it remains challenging to adapt this ability to SVBRDFs, which have more channels, while preserving computational and memory efficiency. Unlike natural images, generating SVBRDFs requires ensuring consistency across different maps (albedo, roughness, normal, etc.).
Extensive research has studied SVBRDF generation from input images (Deschaintre et al., 2018a; Guo et al., 2020; Luo et al., 2024) or text prompts (Vecchio et al., 2024b; He et al., 2023). Most existing methods produce these maps at limited resolutions, typically around $512\times 512$, which restricts their practical use.
Thanks to continuous progress in diffusion models (Rombach et al., 2022; Esser et al., 2024; Lipman et al., 2023), recent work ControlMat (Vecchio et al., 2024a) has pioneered 4K material generation. It generates and merges smaller patches progressively through multiple denoising stages. While the resulting SVBRDFs can reach 4K-resolution, they lacks high-fidelity details. This shortcoming arises because the model is trained from scratch on a limited dataset (Adobe, 2021). Furthermore, this requires considerable computational cost, due to the progressive patch-based generation. The newer (not yet published) work (Vecchio, 2024) has improved the quality of the generated SVBRDFs by finetuning a Stable Diffusion model, but it still requires a dedicated Variational AutoEncoder (VAE) (Kingma and Welling, 2014), trained from scratch on a limited dataset, to ensure consistency across different maps.
In this paper, we introduce HiMat, an approach aimed to generate native high-fidelity 4K SVBRDFs while keeping computational and memory costs acceptable and maintaining cross-map consistency, without training any models from scratch. A key feature of HiMat is its ability to reuse a powerful yet efficient Diffusion Transformer (DiT) backbone, which allows for ultra-resolution generation without sacrificing efficiency. We leverage a linear Diffusion Transformer architecture (Xie et al., 2025a, b), which employs linear attention to improve memory and compute efficiency over the original DiT formulation (Peebles and Xie, 2023), while also offering better scalability than conventional UNet-based models (Podell et al., 2024). Furthermore, we adapt and refine the wavelet-based supervision strategy from Diffusion4K (Zhang et al., 2025b) to enhance the preservation of fine details.
A crucial challenge is maintaining consistency across different SVBRDF maps in a lightweight manner without relying on new VAEs or incurring the overhead of full self-attention (Long et al., 2024; Li et al., 2024, 2025). Our key insight is that, unlike video generation tasks that require full self-attention across many spatially non-aligned frames, multi-map SVBRDF generation typically involves a smaller number of maps, which are closely pixel-aligned. This observation motivates us to avoid costly attention between every pixel of every map, and instead non-destructively add a convolution-based layer across maps that offers stronger local inductive biases and improved efficiency. We term this the CrossStitch module. Notably, this design is compatible with various latent DiT models.
Our results demonstrate that HiMat achieves high-fidelity 4K SVBRDF generation with enhanced quality and reduced computational costs. Our method can generate native 4K-resolution SVBRDFs in just 90 seconds on consumer-level hardware, such as the NVIDIA RTX 4090. Furthermore, our framework can be generalized to other tasks, including intrinsic decomposition. Our main contributions are:
  • A memory-efficient and computationally scalable diffusion model for native ultra-high-resolution SVBRDF generation,
  • A lightweight CrossStitch module that captures localized inter-map dependencies, enabling structural consistency across SVBRDF maps, and can be non-destructively added to existing DiT models,
  • comprehensive experiments demonstrating that our method outperforms existing approaches in both visual quality and efficiency.
Figure 2. HiMat can be utilized for SVBRDF generation and intrinsic decomposition. For SVBRDF generation, HiMat takes a text prompt as input and directly synthesizes $4096\times 4096$ material outputs. For intrinsic decomposition, HiMat encodes the input image into a latent representation, which is then concatenated with noisy latent features and denoised to produce the output. An additional instruction prompt acts as a switcher to control the desired output channel. The core architecture comprises NN stacked DiT blocks, each equipped with a lightweight CrossStitch module to enforce consistency across output maps.
2. Related Work
Materials Generation
Material generation remains a fundamental challenge in computer graphics (Guarnera et al., 2016). Many data-driven approaches (Guo et al., 2020; Sartor and Peers, 2023; Zhou et al., 2022; Luo et al., 2024; Guo et al., 2023) aim to estimate SVBRDFs from single-view images, often leveraging generative models as strong priors. Among these, Guo et al. (2023) propose a patch-based estimation framework capable of estimating ultra-high-resolution SVBRDF. More recently, diffusion models have been adopted for SVBRDF generation. MatFuse (Vecchio et al., 2024b) enables multimodal control and volumetric inpainting, while MaterialPalette (Lopes et al., 2024) and MaterialPicker (Ma et al., 2024) focus on extracting material representations from real-world scenes. To support ultra-high-resolution generation, ControlMat (Zhang et al., 2023), MaterialPalette (Lopes et al., 2024), and StableMaterials (Vecchio, 2024) utilize noise or feature rolling techniques to produce seamless patches, which are subsequently merged and progressively upsampled through multi-stage diffusion process. In contrast, our method enables native 4K SVBRDF generation in a single pass without multi-stage upsampling. This holistic generation approach preserves global spatial coherence, allows for richer and more diverse material patterns, and better captures high-frequency mesostructural detail—particularly critical for close-up rendering and physically accurate surface appearance.
Diffusion Models
Diffusion models have emerged as powerful generative frameworks capable of producing diverse, high-quality images (Yang et al., 2023). To reduce the high computational cost of pixel-space diffusion, Stable Diffusion (Rombach et al., 2022) introduces latent diffusion models (LDM), where the denoising process operates in a compressed latent space via a pre-trained VAE. Building on this foundation, recent advances have improved both model architecture and resolution scaling. Early methods primarily adopt U-Net backbones, while DiT (Peebles and Xie, 2023) demonstrates state-of-the-art performance and serves as the backbone of models such as SD3 (Esser et al., 2024), Flux (Labs, 2024), and Lumina2 (Qin et al., 2025). To enable ultra-high-resolution generation, several works employ training-free fusion (Zhang et al., 2025a; Du et al., 2024) or cascaded upsampling (Pernias et al., 2023; Ren et al., 2024), though these approaches suffer from artifact accumulation. In contrast, PixArt-$\Sigma$ and Sana (Xie et al., 2025a) emphasize generation efficiency but tend to neglect high-frequency details and fine textures, as noted by Diffusion4K (Zhang et al., 2025b). To address this, they introduce wavelet-domain supervision to enhance high-frequency fidelity. Nevertheless, these methods are primarily tailored to the image domain and have not been extended to non-image modalities such as SVBRDF, which demand additional structural consistency across multiple output channels.
Consistent multi-image generation
Consistent multi-image generation is essential for a range of tasks, including identity-preserving synthesis (Huang et al., 2024a; Tewel et al., 2024), video generation (Yang et al., 2024a; Zheng et al., 2024), multi-view generation (Long et al., 2024; Li et al., 2024; Huang et al., 2024b; Jeong et al., 2024), and multi-map material generation (Hong et al., 2024; Zhu et al., 2024; Li et al., 2025). To achieve such consistency, most existing methods leverage attention mechanisms to enforce alignment across images. However, attention mechanisms are known for their quadratic time and memory complexity. While recent alternatives such as linear attention (Katharopoulos et al., 2020) and other state-space models (Dao and Gu, 2024; Yang et al., 2024b), they achieve linear efficiency only when the sequence length is significantly larger than the channel dimension—an assumption that does not hold in multi-map generation tasks where the number of maps is small. For multi-map generation, several works (Liu et al., 2024; Hong et al., 2024; Zhu et al., 2024) adopt multi-branch architectures to maintain structural alignment across outputs. Yet, these approaches introduce substantial parameters and compute overhead, rendering them impractical for ultra-high-resolution generation tasks. In contrast, our method introduces a lightweight CrossStitch module that enables localized inter-map communication with minimal overhead, offering a more practical and scalable solution for consistent high-resolution SVBRDF generation.
3. Preliminaries
3.1. Latent diffusion models
LDM enable high-resolution image synthesis by performing the denoising process in a compact latent space. A VAE, comprising an encoder $\mathcal{E}$ and decoder $\mathcal{D}$, compresses an input image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$ into a latent representation $\boldsymbol{z}_{0}=\mathcal{E}(\boldsymbol{x})\in\mathbb{R}^{\hat{H}\times\hat{W}\times{C}}$. For conditional generation, such as text-to-image synthesis, a prompt $\boldsymbol{y}_{\text{text}}$ is encoded by a text encoder $\tau_{\text{text}}$ into a semantic embedding $\boldsymbol{c}_{\text{text}}$, which serves as conditioning input.
The forward diffusion process corrupts the latent variable $\boldsymbol{z}_{0}$ by gradually adding Gaussian noise:
(1) $\boldsymbol{z}_{t}=\alpha_{t}\cdot\boldsymbol{z}_{0}+\sigma_{t}\cdot{\epsilon},\quad{\epsilon}\sim\mathcal{N}(0,1),$
where $\alpha_{t}$ and $\sigma_{t}$ are schedule-dependent coefficients controlling the signal-to-noise ratio at time step $t$.
In original latent diffusion model (Rombach et al., 2022), the denoising network $\Theta$ is trained to predict the added noise based on:
(2) $\epsilon_{\Theta}(\boldsymbol{z}_{t},t,\boldsymbol{c}_{\text{text}})={\epsilon}.$
Recent variants such as SD3 (Esser et al., 2024) adopt a velocity-based objective via flow matching, where the network predicts the velocity:
(3) $v_{\Theta}(\boldsymbol{z}_{t},t,\boldsymbol{c}_{\text{text}})={\epsilon}-\boldsymbol{z}_{0}.$
The network $\Theta$ is implemented using U-Net architectures or (more recently) DiT, where self-attention enables global spatial interactions within the latent representation, and cross-attention allows conditional information, such as text embeddings, to modulate the denoising process.
3.2. Attention mechanism
Given an input sequence feature $\boldsymbol{s}\in\mathbb{R}^{N\times C}$, where $N$ denotes the sequence length and $C$ represents the channel dimension, the attention operation used in DiT architectures is defined as:
(4) $\text{Attention}({Q},{K},{V})=\text{Softmax}\left(\frac{{Q}{K}^{\top}}{\sqrt{d_{k}}}\right){V},$
where Q,K,V{Q},{K},{V} are the query, key, and value matrices computed via learned linear projections. An additional linear projection is typically applied following the attention computation. While this formulation facilitates expressive global interactions, its quadratic time complexity of $\mathcal{O}(3NC+NC^{2}+N^{2}C)$ and space complexity of $\mathcal{O}(2N^{2}+4NC)$ poses a significant bottleneck.
Some methods (Xie et al., 2025a; Cai et al., 2023) suggest leveraging linear attention mechanisms replace softmax-based version in image-related works:
(5) $\text{LinearAttention}({Q},{K},{V})=\frac{\operatorname{ReLU}(Q)(\operatorname{ReLU}(K)^{\top}V)}{\operatorname{ReLU}(Q)(\operatorname{ReLU}(K)^{\top}\mathbf{1})},$
where $\mathbf{1}$ denotes an all-ones vector used to normalize the attention weights. This alternative offers improved computational efficiency with time complexity of $\mathcal{O}(2NC^{2}+7NC)$ and the space complexity of $\mathcal{O}(N+NC+C^{2})$, which effectively reduces to linear complexity when $N\gg C$.
Figure 3. Detailed design of our DiT block. Our modified architecture incorporates a CrossStitch layer, which enables pixel-localized cross-map communication with minimal overhead. This design enhances structural consistency across multi-image outputs.
4. Method
4.1. Overview
The parameters of SVBRDFs vary, but typically, an SVBRDF of size $H\times W$ consists of base color $\boldsymbol{b}\in\mathbb{R}^{H\times W\times 3}$, normal vector $\boldsymbol{n}\in\mathbb{R}^{H\times W\times 3}$, roughness $\boldsymbol{r}\in\mathbb{R}^{H\times W}$, metallicity $\boldsymbol{m}\in\mathbb{R}^{H\times W}$, and height $\boldsymbol{h}\in\mathbb{R}^{H\times W}$, leading to a set of maps:
(6) $\boldsymbol{I}=\{\boldsymbol{b},\boldsymbol{n},\boldsymbol{r},\boldsymbol{m},\boldsymbol{h}\},$
where $\boldsymbol{I}\in\mathbb{R}^{M\times H\times W\times 3}$. To facilitate efficient processing, we concatenate roughness $\boldsymbol{r}$, metallicity $\boldsymbol{m}$, and height $\boldsymbol{h}$ into a single three-channel image, resulting in $M=3$ in our setting. The key criteria for generation can be summarized as:
  • Fidelity: The model should preserve fine-scale details, ensuring that material textures remain visually plausible and artifact-free even at native 4K-resolution.
  • Efficiency: To scale to such high resolutions, the generation should exhibit memory and computational efficiency, avoiding excessive parameter overhead and runtime costs.
  • Consistency: All SVBRDF channels should remain structurally and semantically aligned to maintain photometric and semantic coherence across base color, normals, roughness, metallicity, and height.
To achieve these goals, we design HiMat, which is built on a high-resolution linear-attention DiT (Xie et al., 2025a) as its backbone. By substituting the original quadratic self-attention mechanism with linear attention, the linear DiT improves both memory and computational efficiency, which is crucial for synthesizing 4K-resolution outputs.
At the core of HiMat is CrossStitch, a lightweight $1\times 1$ convolution module designed to ensure precise structural alignment across SVBRDF channels. To capture and enhance intricate high-frequency details, we also employ a wavelet-based fine-tuning strategy that leverages the power of the stationary wavelet transform.
Unlike prior works (Vecchio et al., 2024b, a; Vecchio, 2024) that trained material autoencoders from scratch, we leverage a pre-trained DC-AE (Chen et al., 2025). This decision is motivated by two main reasons. First, due to the scarcity of material data, training from scratch risks learning a suboptimal latent representation, which can result in lost details. Second, prior work demonstrated only a limited $8\times$ compression ratio, while the pre-trained DC-AE allows for a higher compression ratio of $32\times$, facilitating more efficient 4K material generation. Specifically, given the input material maps $\boldsymbol{I}$, we encode them into latent space as $\boldsymbol{z}_{0}=\mathcal{E}(\boldsymbol{I})\in\mathbb{R}^{M\times\hat{H}\times\hat{W}\times C}$, where $\hat{H}=\frac{H}{32}$, $\hat{W}=\frac{W}{32}$, and $C=32$.
Figure 4. Comparison between HiMat and StableMaterials (Vecchio, 2024) for text-to-material generation. Note that StableMaterials is not yet published, and currently supports only 512 × 512 resolution. Our results tend to have better detail and feature scale. See Tab. 1 for quantitative metrics, and the supplementary for many more results.
4.2. CrossStitch: an inter-map correlation-aware module
To ensure consistency across SVBRDF maps, the key lies in facilitating effective information exchange among them. Most existing approaches (Li et al., 2025; Long et al., 2024; Li et al., 2024; Ma et al., 2024) employ full 3D self-attention, enabling joint modeling across spatial and materials dimensions at the cost of significant computational and memory costs due to its quadratic complexity (Zheng et al., 2024). While the decomposition of attention into separate spatial and cross-map branches can reduce overhead, it still remains suboptimal for SVBRDF generation. Our key observation is that, unlike video generation, which necessitates long-range temporal context, SVBRDF generation only involves a limited number of pixel-aligned maps. Therefore, self-attention across maps is unnecessary.
To this end, we propose CrossStitch, a lightweight design that localizes inter-map correlations efficiently, as shown in Fig. 3. Specifically, given a latent map $\boldsymbol{f}\in\mathbb{R}^{M\times\hat{H}\times\hat{W}\times\hat{C}}$ after self-attention, where $\hat{C}$ denotes the attention feature dimension, CrossStitch performs local information exchange along the map dimension by applying a mapping along the material map dimension $M$, and restoring the original layout (using einops (Rogozhnikov, 2022) notation):
[…]
$\boldsymbol{f}$ $\leftarrow\mathrm{CrossStitch}(\boldsymbol{f})$
[…]
where we use m h w c to denote the symbolic dimensions $M$, $\hat{H}$, $\hat{W}$, and $\hat{C}$ in the rearrange string.
The CrossStitch consists of depthwise separable convolutions-a depthwise convolution with kernel size $M$, followed by a pointwise $1\times 1$ convolution for channel mixing. Additionally, a parallel branch applies average pooling across the map dimension followed by a $1\times 1$ convolution, enabling the module to aggregate global structural context across maps. Intuitively, this branch computes the mean semantic information across maps, facilitating inter-map alignment. We zero-initialize all layers of CrossStitch to ensure that the initial output does not disrupt the original feature space. This architectural choice allows the model to build upon the established priors.
4.3. Wavelet-based fine-tuning
For 4K-resolution SVBRDF maps, our goal is not merely to increase spatial resolution, but to enhance content fidelity: preserving and generating richer details. To do that, Diffusion4K (Zhang et al., 2025b) explicitly supervises the diffusion model in the frequency domain based on Eqn. (3):
(8) $\mathcal{L}_{{DWT}}=\mathbb{E}\left[\lambda\|\mathcal{W}_{DWT}(\boldsymbol{v}_{\Theta}(\boldsymbol{z}_{t},t))-\mathcal{W}_{DWT}(\epsilon-\boldsymbol{z}_{0})\|^{2}\right],$
where $\lambda$ is the loss weight, and $\mathcal{W}_{DWT}(\cdot)$ denotes Discrete Wavelet Transform (DWT), which decomposes the signal into four orthogonal subbands: a low-frequency approximation LL and three high-frequency components LH, HL, HH. However, the inherent downsampling in DWT reduces the spatial resolution of each subband by half, limiting its capacity to effectively supervise high-resolution structure.
To overcome these limitations, we adopt a Stationary Wavelet Transform (SWT)-based loss, which has been shown effective in super-resolution tasks (Korkmaz et al., 2024; Korkmaz and Tekalp, 2024). It avoids downsampling and retains full spatial resolution across subbands, making it well-suited for high-resolution generation. The final loss is formulated as:
(9) $\mathcal{L}_{{SWT}}=\mathbb{E}\left[\sum_{i}\lambda_{i}\left\|\mathcal{W}_{{SWT}}(\boldsymbol{v}_{\Theta}(\boldsymbol{z}_{t},t))_{i}-\mathcal{W}_{{SWT}}(\epsilon-\boldsymbol{z}_{0})_{i}\right\|^{2}\right],$
where $i\in\{\text{LL},\text{LH},\text{HL},\text{HH}\}$, $\mathcal{W}_{{SWT}}(\cdot)_{i}$ denotes the $i$-th SWT subband, and $\lambda_{i}$ is its associated loss weight. As natural signals concentrate most energy in the low-frequency LL subband, gradients from high-frequency bands are typically underemphasized. To compensate, we use higher weights for high-frequency components: $\lambda_{\text{LL}}=0.5$, $\lambda_{\text{LH}}=\lambda_{\text{HL}}=2$, and $\lambda_{\text{HH}}=2.5$, using the Symlet-19 wavelet basis.
5. Implementation & results
Table 1. Quantitative comparison for SVBRDF generation with StableMaterials (Vecchio, 2024).
StableMaterials Ours
CLIPScore $\uparrow$ 24.14 24.34
Aesthetics $\uparrow$ 3.82 4.43
HPS $\uparrow$ 0.19 0.20
GLCM Score $\uparrow$ 0.54 1.34
5.1. SVBRDF generation
Dataset
We train our model on the combined datasets of MatSynth (Vecchio and Deschaintre, 2024) and Deschaintre (2018b), for a total of 6,198 unique Physically Based Rendering (PBR) materials. Also, we perform data augmentations like horizontal flipping, vertical flipping and random rotation.
Technical details
We initialize training from a pre-trained Sana-1024px checkpoint with 20 DiT blocks, 1.6B parameters. To effectively learn high-frequency content while maintaining stability, we adopt a progressive strategy, gradually increasing the training resolution from $1024\times 1024$ to $2048\times 2048$, and finally to $4096\times 4096$. We use noise-rolling (Vecchio et al., 2024a) to enable tileable material generation.
Comparison
We compare our method against StableMaterials (Vecchio, 2024), which has demonstrated superior results over MatFuse (Vecchio et al., 2024b) and ControlMat (Vecchio et al., 2024a). It is worth noting that StableMaterials is not yet published, and currently supports only $512\times 512$ resolution. To assess image quality, we conduct comprehensive evaluations using standard perceptual metrics, including CLIPScore (Hessel et al., 2021), Aesthetics (Schuhmann et al., 2022), and Human Preference Score (HPS) (Wu et al., 2023), providing intuitive measures of visual quality and prompt alignment. We additionally report the GLCM Score (Zhang et al., 2025b), which evaluates texture richness in ultra-high-resolution images. All metrics are computed on rendered images, and the full list of 202 text prompts is provided in the supplemental materials.
Results
As shown in Fig. 4 and Fig. 7, our results closely align with the input prompts, generating diverse and realistic materials even for out-of-distribution concepts. Quantitative results in Tab. 1 further confirm that our method yields higher-quality outputs with richer details, reflected by improved Aesthetics scores, and GLCM scores. Additional qualitative examples are provided in the supplemental materials.
Figure 5. Visual comparison with 4K generation methods. We compare our approach against Diffusion4K (Zhang et al., 2025b) and Sana-4K (Xie et al., 2025a).
Comparison to 4K text-to-image methods
To further contextualize our performance, we compare HiMat with existing 4K text-to-image generation models: Diffusion4K (Zhang et al., 2025b) and Sana-4K (Xie et al., 2025a). Diffusion4K fine-tunes Flux-12B (Labs, 2024) using wavelet-based supervision, while Sana-4K is designed for fast 4K generation via architectural optimizations. Note that this is a very strict test for the realism of our method, as it has a much harder task of generating SVBRDF maps which are then rendered under environment lighting, while the other methods are standard RGB image generators. As shown in Fig. 5, Sana-4K lacks coherent structural details and exhibits over-saturation, likely due to its reliance on super-resolved training data rather than native 4K samples. Diffusion4K struggles with texture generation and incurs high inference costs with 12B model. In contrast, HiMat achieves visually rich and structurally faithful 4K rendered outputs, demonstrating superior high-frequency detail preservation and semantic fidelity.
Native 4K vs. super-resolution
We investigate the difference between native 4K generation and super-resolution-based approaches. To this end, we first generate $1024\times 1024$ outputs using our 1024-resolution model, and then apply SUPIR (Yu et al., 2024)—a state-of-the-art super-resolution method—to upscale the results to $4096\times 4096$ for each map. As shown in Fig. 6, our native 4K generation achieves better detail preservation. While SUPIR enhances image sharpness, it introduces color bias.
Figure 6. Visual comparison between our native 4K SVBRDF generation and the super-resolution method SPUIR (Yu et al., 2024). SPUIR introduces noticeable color bias.
Table 2. Quantitative comparison on Hypersim (Roberts et al., 2021) test dataset.
Ours w/o CrossStitch RGB→X (2024)
[…]
Normal 16.52 15.02 19.8
Irradiance 16.34 15.34 14.1
Mean 15.61 14.25 17.1
[…]
Normal 0.26 0.33 0.26
Irradiance 0.26 0.33 0.22
Mean 0.27 0.35 0.19
5.2. Intrinsic decomposition
Our CrossStitch module is compatible with a wide range of latent diffusion models and transferable across different tasks. To demonstrate its effectiveness beyond material generation, we additionally conduct experiments on intrinsic image decomposition using SD3.5-Medium (Esser et al., 2024).
Table 3. Comparison of computational cost during inference. All results are measured on a consumer level RTX 4090D GPU. We report both the absolute results and the relative ratio (normalized by CrossStitch variant, denoted as 1.00) for variants using attention and linear attention.
[…]
1024×1024 2048×2048 4096×4096 1024×1024 2048×2048 4096×4096 1024×1024 2048×2048 4096×4096
Attention 2.01 / 1.14 11.20 / 1.21 43.68 / 1.22 173.58 / 1.22 10.89 / 1.05 12.76 / 1.01 20.44 / 1.03 0.39 / 1.30 1.23 / 1.29 5.13 / 1.28
Lin. ttn. 2.01 / 1.14 11.22 / 1.21 43.75 / 1.22 173.88 / 1.22 11.59 / 1.12 15.80 / 1.25 OOM 0.43 / 1.43 1.30 / 1.37 -
Ours (CrossStitch) 1.76 / 1.00 9.25 / 1.00 35.87 / 1.00 142.36 / 1.00 10.39 / 1.00 12.63 / 1.00 19.93 / 1.00 0.30 / 1.00 0.95 / 1.00 4.01 / 1.00
Dataset
We conduct our experiments on the Hypersim dataset (Roberts et al., 2021), which contains 74,619 images of indoor scenes, each annotated with albedo, normal, and irradiance maps. We use the official training split for model training and evaluate performance on the test split. For consistency and fair comparison with RGB→X (Zeng et al., 2024), all data is resized to $512\times 512$ resolution.
Comparison
[…]
Results
[…]
5.3. Ablation studies
Benefits of CrossStitch.
We evaluate the efficiency of our CrossStitch-based architecture by comparing it against standard attention and linear attention baselines across varying resolutions based on Sana, as summarized in Tab. 3. We replace the CrossStitch module with standard or linear attention layers, and measure their performance respectively. We report the number of parameters, forward FLOPs, peak memory usage, and inference time per step, with all experiments conducted on a consumer-level RTX 4090D GPU. Our Cross-Stitch module achieves consistent reductions in computational overhead across all scales. At $1024\times 1024$ and $2048\times 2048$ resolutions, it delivers up to $22\%$ fewer FLOPs and $25\%$ lower memory consumption compared to linear attention, while also yielding faster inference. At $4096\times 4096$ resolution, standard attention and linear attention both incur prohibitive cost, with the latter even leading to out-of-memory failure.
Table 4. Quantitative comparison of wavelet-based supervision strategies at $1024\times 1024$ resolution.
Naive DWT-based Loss SWT-based Loss
GLCM Score $\uparrow$ 0.61 0.64 0.69
Effect of wavelet-based loss
We ablate the impact of our wavelet-based fine-tuning strategy by comparing models trained with and without the proposed SWT-based supervision at a resolution of $1024\times 1024$. We additionally compare against the DWT-based method introduced in Diffusion4K (Zhang et al., 2025b). As shown in Tab. 4, incorporating wavelet-domain supervision substantially improves the preservation of high-frequency geometric features, as reflected in higher GLCM scores. Moreover, our SWT-based loss proves more effective than DWT by avoiding downsampling, thus providing finer spatial fidelity and stronger supervision for detailed structure.
Figure 7. Text prompting for 4K SVBRDF generation. Zoom-in views are provided to highlight details. We present a diverse set of materials generated from text prompts, demonstrating the versatility of our method. Additional results are provided in the supplemental material.
Figure 8. Qualitative comparison against previous method (Zeng et al., 2024) and our variant without CrossStitch on the Hypersim testset (Roberts et al., 2021).
5.4. Discussion and limitations
Our method still has several limitations. The most significant being the challenge of finetuning under limited data. Unlike state-of-the-art image generation models that benefit from large-scale RGB image datasets, native 4K SVBRDF datasets (especially openly available ones) remain scarce. Although prior efforts (Ma et al., 2023; Deschaintre et al., 2023; Chen et al., 2024) have introduced material datasets, they often face constraints such as limited public availability, small data volume, lack of true 4K resolution, and insufficiently rich textual annotations. Additionally, our model comprises only 1.8B parameters, making it less directly comparable to large-scale architectures such as Flux-12B (Labs, 2024) and HiDream-I1-17B (HiDream.ai, 2025). Nonetheless, our design remains scalable and easily to extend with methods introduced by Sana1.5 (Xie et al., 2025b). Finally, while our current implementation supports only text conditioning for SVBRDF generation, it is compatible with extensions to support image and other conditional modalities like ControlMat (Vecchio et al., 2024a), as evidenced by the intrinsic decomposition task. We leave this for future work.
6. Conclusion
We propose HiMat, an efficient DiT-based framework for native 4K SVBRDF generation. By finetuning a linear-attention DiT with lightweight design choices, HiMat preserves consistency, fidelity and efficiency of the original model. A key contribution of HiMat is CrossStitch, a non-destructive module compatible with most DiT image generation models, designed to replace computationally expensive attention across all pixels of all maps, while preserving cross-map alignment. This module leverages localized inductive bias to capture structural dependencies across SVBRDF channels with minimal parameter cost. We further enhance visual fidelity through frequency-aware supervision using stationary wavelet transforms, which improves the generation of microstructure-level detail at 4K resolution. Our experiments demonstrate that HiMat delivers superior visual quality and structural consistency while still delivering 4K resolution outputs on a consumer-level GPU. We believe HiMat offers a strong foundation for scalable, high-quality SVBRDF generation and opens the door to broader adoption of efficient DiT-based pipelines in digital content creation.